# A Study of Teleoperation Methods in a Simulated Virtual Eye Surgery Environment

### v1

@@ Abstract
This paper examines the performance of Inside and Outside Control modes at various scaling factors in a simulated vitreoretinal surgical setting. The IRISS teleoperated surgical system’s console (cockpit) was adapted to project a simulated microscope view of an intraocular setup to a virtual reality (VR) headset. Five experienced vitreoretinal surgeons and five engineers with no surgical experience used the system to perform tasks common to vitreoretinal surgery. Experimental results indicate that Inside Control methods at higher scaling factors (20 or 30) achieved the best performance overall, though the optimal scaling factor may vary by task and complexity. Optimizing control methods and scaling factors could lead to improvements in surgical efficiency and accuracy, as well as minimize risks in future robotic-assisted intraocular procedures.
Keywords: Virtual Reality Simulation, Teleoperation, Ophthalmology, Inside Control, Outside Control, Scaling Factor
Figure 1: Experimental Teleoperation Setup (a) Master Manipulator and Cockpit (controlled by the surgeon). (b) Closeup view of Master Manipulator Joystick (c) Closeup view of Joystick Gripper Mechanism
@@ I Introduction
Vitreoretinal microsurgery is among the most challenging surgical procedures in ophthalmology, requiring high precision in extremely constrained spaces. The delicate nature of the procedure stems from the need to manipulate tissues on the micron scale, as even the largest retinal veins measure only around $150\mu m$ in diameter [1] [2]. This intricate work demands exceptional dexterity and control, as any unintended motion can result in irreversible damage to retinal structures. Additionally, the surgical environment poses unique challenges, such as reduced spatial resolution and depth perception due to the reliance on microscopic visualization [3]. Compounding these issues are physiological factors, including hand tremors and a limited tactile response, which further elevate the difficulty of achieving consistent surgical outcomes [4, 5, 6]. The aforementioned challenges of vitreoretinal surgery make it an optimal candidate for robotic intervention and success in addressing many of these issues has been demonstrated in the past with several robotic surgical platforms. In particular, minimization of hand tremor and improvements to surgical precision have been noted as benefits [7, 8, 9]. Current approaches to surgery range from fully manual to fully automated robotic systems, with a spectrum of assistance levels in between. Teleoperated surgery lies within this spectrum. In this form of robot assisted surgery, the surgeon uses a microscope or a digital heads-up display (HUD) as visual feedback while controlling the robotic system via a spatially representative control system, like joysticks [10].
@@
Another system, designed more specifically for intraocular surgery, is the Preceyes Surgical System (Preceyes BV, Eindhoven, the Netherlands), which has a remote center of motion and a joystick that controls the robot’s movements outside of the eye, providing inverse mapping [21, 22]. These contradictions in literature and robotic design choices can be attributed to inconsistent experimental setups, choice of test participants, and chosen surgical procedures. Additionally, no existing research specifically addresses the unique challenges of intraocular surgery in their comparisons. As such, further study is required to determine the ideal control mode for teleoperated ophthalmic surgery.
@@ III Results and Discussion
Figure 7: Performance for Task 1
@@
Figure 8: Performance for Task 2
@@
Figure 9: Performance for Task 3
@@
Figure 10: Performance for Task 4

### v2

# Control Modes of Teleoperated Surgical Robotic System’s Tools in Ophthalmic Surgery

@@ Abstract
The introduction of a teleoperated surgical robotic system designed for minimally invasive procedures enables the emulation of two distinct control modes through a dedicated input device of the surgical console: (1) Inside Control Mode, which emulates tool manipulation near the distal end (i.e., as if the surgeon was holding the tip of the instrument inside the patient’s body), and (2) Outside Control Mode, which emulates manipulation near the proximal end (i.e., as if the surgeon was holding the tool externally). The overarching aim of this reported research is to study and compare the surgeon’s performance utilizing these two control modes of operation along with various scaling factors in a simulated vitreoretinal surgical setting. The console of Intraocular Robotic Interventional Surgical System (IRISS) was utilized but the surgical robot itself and the human eye anatomy was simulated by a virtual environment (VR) projected microscope view of an intraocular setup to a VR headset. Five experienced vitreoretinal surgeons and five subjects with no surgical experience used the system to perform fundamental tool/tissue tasks common to vitreoretinal surgery including: (1) touch and reset; (2) grasp and drop; (3) inject; (4) circular tracking. The results indicate that Inside Control outperforms Outside Control across multiple tasks and performance metrics. Higher scaling factors (20 and 30) generally provided better performance, particularly for reducing trajectory errors and tissue damage. This improvement suggests that larger scaling factors enable more precise control, making them the preferred option for fine manipulation tasks. However, task completion time was not consistently reduced across all conditions, indicating that surgeons may need to balance speed and accuracy/precision based on specific surgical requirements. By optimizing control dynamics and user interface, robotic teleoperation has the potential to reduce complications, enhance surgical dexterity, and expand the accessibility of high-precision procedures to a broader range of practitioners.
Keywords: Virtual Reality Simulation, Teleoperation, Ophthalmology, Scaling Factor, Control Mode, Surgical Robotics, Ophthalmic Surgery, Ophthalmic Surgical Robotic System
Figure 1: Experimental Teleoperation Setup (a) Master Manipulator and Cockpit (controlled by the surgeon). (b) Closeup view of Master Manipulator input device (c) Closeup view of input device Gripper Mechanism
@@ I Introduction
In Minimally Invasive Surgery (MIS), surgical instruments are introduced into the body through small ports established at the skin surface or, in the case of ophthalmic procedures, through specific ocular tissues such as the sclera, cornea, or conjunctiva. The establishment of these ports fundamentally alters the surgeon’s interaction with the instrument. Unlike open surgery, where the surgeon may manipulate the tool from any position along its shaft—including proximally or distally—MIS confines the surgeon’s interaction to the proximal end of the tool, which remains external to the patient’s body, while the distal end performs the intervention through the fixed port.
The presence of a port introduces several critical constraints: (a) It reduces the number of degrees of freedom (DOF) available for tool manipulation from six to four (excluding grasping), unless wristed instruments are employed; (b) It transforms tool motion from Cartesian space (linear and rotational movement in 3D) into a spherical motion space, pivoting around the port; (c) It inverts the relationship between handle and tip motion, such that movement of the handle in one direction results in motion of the tool tip in the opposite direction inside the body.
To address these limitations and enhance surgical control, the use of teleoperated robotic systems in MIS enables the emulation of two distinct control modes through a dedicated input device at the surgical console: (1) Inside Control Mode, which emulates tool manipulation near the distal end (i.e., as if the surgeon was holding the tip of the instrument inside the patient’s body), and (2) Outside Control Mode, which emulates manipulation near the proximal end (i.e., as if the surgeon was holding the tool externally).
Vitreoretinal microsurgery is among the most challenging surgical procedures in ophthalmology, requiring high precision in extremely constrained spaces. The delicate nature of the procedure stems from the need to manipulate tissues on the micron scale, as even the largest retinal veins measure only around $150\mu m$ in diameter [1] [2]. This intricate work demands exceptional dexterity and control, as any unintended motion can result in irreversible damage to retinal structures. Additionally, the surgical environment poses unique challenges, such as reduced spatial resolution and depth perception due to the reliance on microscopic visualization [3]. Compounding these issues are physiological factors, including hand tremors and a limited tactile response, which further elevate the difficulty of achieving consistent surgical outcomes [4, 5, 6]. The aforementioned challenges of vitreoretinal surgery make it an optimal candidate for robotic intervention and success in addressing many of these issues has been demonstrated in the past with several robotic surgical platforms. In particular, minimization of hand tremor and improvements to surgical precision have been noted as benefits [7, 8, 9]. Current approaches to surgery range from fully manual to fully automated robotic systems, with a spectrum of assistance levels in between. Teleoperated surgery lies within this spectrum. In this form of robot assisted surgery, the surgeon uses a microscope or a digital heads mounted display (HMD) as visual feedback while controlling the robotic system via a spatially representative control system, like input devices [10].
@@
Another system, designed more specifically for intraocular surgery, is the Preceyes Surgical System (Preceyes BV, Eindhoven, the Netherlands), which has a remote center of motion and a input device that controls the robot’s movements outside of the eye, providing inverse mapping [21, 22]. These contradictions in literature and robotic design choices can be attributed to inconsistent experimental setups, choice of test participants, and chosen surgical procedures. Additionally, no existing research specifically addresses the unique challenges of intraocular surgery in their comparisons. As such, further study is required to determine the ideal control mode for teleoperated ophthalmic surgery.
@@ III Results
Figure 7: Performance for Task 1: Touch and reset
@@
Figure 8: Performance for Task 2: Grasp and drop
@@
Figure 9: Performance for Task 3: njection
@@
Figure 10: Performance for Task 4: Circular tracking
@@ IV Conclusions
There are two key ways in which minimally invasive (MIS) ophthalmic surgery differs from MIS procedures in general when utilizing teleoperated surgical robotic systems. These differences have important implications for the control mode paradigms. (1) Tool Control Responsibility and Workflow: In ophthalmic surgery, the surgeon maintains continuous control of the instrument via the teleoperation console before, during, and after insertion into the eye. The surgical assistant (nurse) mounts the instrument onto the robotic arm, but the surgeon actively controls the tool throughout its entire trajectory, including the critical phase of insertion through the ocular port (e.g., the sclera or cornea). This contrasts with general MIS procedures, where the instrument is typically inserted into the patient by the assistant until the tip is visible in the endoscopic field of view. Only then does the surgeon assume control via the input device. As a result, ophthalmic surgery requires a more comprehensive control strategy, and often relies on an extended visual field, such as a wide-angle digital microscope view that includes both intraocular and extraocular portions of the instrument. (2) Spatial Relationship Between Port and Surgical Site: In ophthalmic surgery, the surgical site is located very close to the port of entry, and both the intraocular and extraocular portions of the tool are often visible in the surgical microscope’s field of view. In contrast, general MIS typically involves longer instruments, where the operative site is located deeper within the body and far from the entry port. The external portion of the tool remains outside the camera’s view, as the endoscope is focused tightly on the internal surgical field. This spatial difference affects how the surgeon perceives tool motion and may influence the control mode preferences. It is important to note that these unique characteristics of ophthalmic MIS were not directly investigated in the present study. However, they may significantly influence the choice of control mode—specifically between inside and outside Control Modes. For instance, cataract surgery, where the anatomical target is very close to the incision, may favor one control mode, whereas vitreoretinal procedures (simulated in this study), where the target is located deeper within the eye, may benefit from a different mode.